OpenStudy (anonymous):

What is the solution to the equation 3^1 + 2x = 2

OpenStudy (anonymous):

3^1=3 3+2x=2 2x=2-3 2x=-1 x=-1/2

OpenStudy (asnaseer):

I /think/ your question is:\[3^{1+2x}=2\]if so, you just need to take logs of both sides to get:\[\log(3^{1+2x})=\log(2)\]\[(1+2x)\log(3)=\log(2)\]\[1+2x=\frac{\log(2)}{\log(3)}\]\[2x=\frac{\log(2)}{\log(3)}-1\]\[x=\frac{1}{2}\left(\frac{\log(2)}{\log(3)}-1\right)\]